# access point PTP Precision Time Protocol

### v1

@@ Abstract
The proliferation of 5G technology necessitates advanced network management strategies to ensure optimal performance and reliability. Digital Twin (DT)s have emerged as a promising paradigm for modeling and simulating complex systems like the 5G Radio Access Network (RAN). In this paper, we present TwinRAN, a DT of the 5G RAN built leveraging the Azure DT platform. TwinRAN is built on top of the Open RAN (O-RAN) architecture and is agnostic to the vendor of the underlying equipment. We demonstrate three applications using TwinRAN and evaluate the required resources and their performance for a network with  800 users and 8 gNBs.
We first evaluate the performance and limitations of the Azure DT platform, measuring the latency under different conditions. The results from this evaluation allow us to optimize TwinRAN to the DT platform it uses. Then, we present the system’s architectural design, emphasizing its components and interactions. We propose that two types of twin graphs be simultaneously maintained on the cloud. The first one is for intercell operations, keeping a broad overview of all the cells in the network. The second twin graph where each cell is spawned in a separate Azure DT instance for more granular operation and monitoring of intracell tasks. We evaluate the performance and operating costs of TwinRAN for each of the three applications. The TwinRAN DT in the cloud can keep track of its physical twin within a few hundred milliseconds, extending its utility to many 5G network management tasks - some of which are shown in this paper. The novel framework for building and maintaining a DT of the 5G RAN presented in this paper offers network operators enhanced capabilities, empowering efficient deployments and management.
@@ I Introduction
DTs, virtual replicas of physical systems, offer a promising approach to addressing these challenges [5]. Operators can simulate, monitor, and optimize network performance in real-time by creating a DT of the 5G RAN, enabling predictive management, efficient resource allocation, and rapid troubleshooting [6, 7]. DTs are being leveraged in many other industries, such as manufacturing, logistics, construction, and energy distribution [8]. At the forefront of the DT technology is the Microsoft Azure DT[9]. It is a Platform as a service (PaaS) to create, maintain, and interact with DTs on the cloud. The Azure cloud, where the DTs are hosted, is a production-grade cloud platform guaranteeing scalability, availability, and robustness. Moreover, other Azure cloud storage, analytics, and visualization services can be seamlessly connected to the DTs. Therefore, Azure DT presents a compelling opportunity for twinning the 5G RAN. Hence, we first measure the real-time performance of Azure DT under different conditions to make optimized design decisions for TwinRAN. The presented results from these measurements are a supplementary contribution to TwinRAN. To the best of our knowledge, this paper is the first to present a deep study of the behavior of Azure DT.
@@
This paper proposes a novel method and architecture for creating and managing a DT of the 5G RAN on the Azure cloud platform. Our approach leverages Azure DT’s standardized modeling, data integration, and analytical capabilities to provide a RAN. TwinRAN has three main features that underscore its utility: Non-Invasive: It doesn’t require changes in the source code or hardware in the 5G RAN. By plugging the xApp on top of 5G RAN architectures, it builds seamlessly over the existing 5G components. Vendor Agnostic: It can work with all equipment and software vendors as long as they are O-RAN compliant. Multipurpose: The same DT can be simultaneously used for different applications. Any application can connect to the DT to stream real-time updates or load historical data.
The rest of the paper is organized as follows: the next section provides the background on Azure DTs and the O-RAN system. Following the background, we mention some related work and differentiate our contributions. In section IV, we evaluate the behavior of Azure DT. The learnings from this section are then used to present the architecture of TwinRAN in Section V. This section also elaborates on the Digital Twins Definition Language (DTDL) models and the xApp design. Section 6 addresses the implementation of TwinRAN and introduces the three use cases, and Section VII evaluates the performance and required resources for TwinRAN, mainly focusing on pricing. Finally, in Section VIII, we conclude, presenting the limitations and point toward out future work.
@@ II-A4 Message Types
O-RAN defines several message types for communication between its components. Indication: Provides status updates and telemetry data from RAN elements to the Near-RT RIC. The two types of indication messages are report and insert. While the first gives periodic updates on subscription, the other is triggered asynchronously to notify the xApp of a specific event at the RAN. Policy: Convey policies from the Near-RT RIC to guide the behavior of RAN elements. The passing of such a policy function is reported to the TwinRAN instance by the TwinRAN xApp. Control: Used for configuring and managing RAN elements. While TwinRAN does not use control messages, we show an example in Section 6 how they can be used in conjunction with TwinRAN to perform a handover.
@@ II-B1 Models
Models in Azure DT define the structure and properties of DTs. These models are created using the DTDL, specifying entities’ characteristics, behavior, and capabilities within the system. DTDL, is JSON-like and comprises a set of metamodel classes  [20]. Defining clear and consistent models ensures a standardized representation of the physical entities and their attributes [21, 19]. The main metamodel classes are:
Interface: Defines the blueprint for a digital twin, outlining its structure, properties, and capabilities.
Property: Describes attributes of a digital twin, holding specific static values. For example, the serial number of the thermometer.
Relationship: Details the connections between digital twins, specifying how they interact and relate to each other—for example, the room and the thermometer and thermostat.
Component: Represents a sub-entity of a digital twin by linking to an interface. It allows for modularity in defining models—for example, the window in a room. DTDL defines more classes, such as command and telemetry, which do not have a unique advantage in DTs [22] They are mainly useful for other Azure IoT services.
@@ III-B Digital Twin Platforms
Eclipse Ditto[28]An open-source framework part of the Eclipse IoT Working Group. It provides the tools to create, manage, and interact with digital twins. Similar to Azure DT, it has a JSON-based twin model defining language. However, twin instances are not maintained as knowledge graphs as in Azure DTs. It can be self-hosted and connected to various MQTT and HTTP protocol endpoints.
MCX:  [44] An open-source model description and updating method via MQTT protocol. It connects to simulation frameworks such as MATLAB, Ansys, or custom Python classes. It facilitates connecting simulation frameworks with real-time data.
AWS IoT TwinMaker[45] AWS IoT TwinMaker is a cloud-native platform rivaling the Azure DT platform in its capabilities and features. It connects to other AWS services for monitoring, analytics, and data ingestion. The models are defined in a JSON-like language, and the twin instances are maintained in a knowledge graph. Many web tools like Grafana dashboards are available for twin visualization and monitoring.
@@ IV Evaluating Azure DT
To ascertain whether any two distributions in our measurements are the same, we employ the Kolmogorov-Smirnov test with a p-value threshold of 0.05. The time of the day did not make any difference to any of the three values mentioned above. Similarly, updating or querying multiple twins on the same Azure DT instance did not increase either of the three values. When the model and update sizes were 100, we measured no difference between the first and last parameters’ lag. This indicates that the parameters on a twin are either not updated sequentially but rather parallelly, or the time to update all the parameters is very small compared to the overall lag and service time.
@@ V-A Intercell and Intracell Instances
The limits placed by Azure DT provide us with considerable headroom on the message size. Let the maximum number of twins in an instance be $Y\in\mathbb{N}^{+}$. If every twin in the instance is updated once every $R\in\mathbb{R}^{+}$ seconds, then the Azure DT constraint is $R\geq 0.1$, and we get $Y=1000\cdot R$. 5G RAN systems are dynamic, and the maximum frame and slot length are 10 and 1 ms, respectively. Therefore, to maintain the fidelity of the DT in the cloud, the update rate must be as high as possible. Updating all the twins at the maximum rate of $R=0.1$ seconds, we can have a maximum of 100 twins per instance.
From the perspective of a 5G RAN operator, the tasks can be divided into two main categories: intracell and intercell. Intracell operation tasks include efficient scheduling of users based on their required Quality of Service (QoS), minimizing operating costs, network security, etc. Intercell operations are handover management, interference mitigation, etc. Therefore, we propose two twin-graphs simultaneously maintained in the cloud: TwinRAN-cell and TwinRAN-multi. TwinRAN-multi keeps a broad overview of the entire network. Each cell is defined in one DTDL model and neighboring cells are connected with relationships. Each UE is one property of the gNB, with SINR as its value. TwinRAN-cell spawns a new instance for every gNB. Here, the gNB and UE are two different DTDL models with a relationship between them if the UE is connected. We present the DTDL models in detail in the following subsection. A large model of a cell with detailed models of the gNB and UE as components would allow us to simultaneously manage 100 cells, each updating 10 times a second. However, a cell containing 100s of UEs would make the update sizes large and increase the twins’ lag, service time, and query time. Moreover, this would impact the system’s scalability as an additional parameter in the UE model and would severely impact the overall lag. We can keep these values low by employing a disaggregated approach that simultaneously maintains TwinRAN-cell and TwinRAN-multi instances.
@@ V-B DTDL Models
As mentioned in Section II-B, models defined by DTDL provide a structure to each twin. TwinRAN defines three models: the cell, User Equipment (UE), and the gNB. The cell is used in TwinRAN-multi. The input to our modeling step is the information we can obtain via different O-RAN messages mentioned in Section II-A. The information is divided into Information Elements (IE)s, which are structured data units used within the O-RAN protocol messages to convey specific parameters. While the amount of information is substantial, and thus, one can make very detailed models, we make the smallest possible model useful for all the three use cases mentioned in Section 6.
Table I shows the two DTDL interfaces for the TwinRAN-multi instance. The gNB is already initialized with a given number of UEs regardless of their connection. This is because changing a model in the live execution environment requires more than 1 second. Neighboring gNBs are connected with a relationship whose value is the received power from that gNB. The properties of the UE that are important to demonstrate our use cases are its location and received power.
@@
Tx Power float 20-50 DL MCS int 0-15
Relationship int UE DL BLER float 0.0-1.0
Connected UEs int 0-99 DL MCS int 0-15
@@ V-C TwinRAN xApp
The TwinRAN xApp updates the DT on the cloud and manages the instances, creating twins and maintaining relationships. The information obtained from the E2 Node is provided by different RAN functions [18]. TwinRAN subscribes to the required RAN functions via the Near-RT RIC as shown in Figure 1. For periodic indication messages(called as report II-A, the subscription request also contains a parameter called granularity. This parameter represents the frequency at which the report messages should be sent by the E2 Node. The gNB in the E2 setup request provides a cell ID and PLMN information used to spawn a new cell in the TwinRAN-muli instance. Meanwhile, another instance is created in the Azure DT cloud, where the gNB twin is created. One expects this to happen rarely, as cells do not frequently appear and disappear from the operator’s network.
@@ VI-A Use Case I: Data Warehouse for Machine Learning
A data warehouse is a centralized repository that stores large volumes of structured data from multiple sources. Access to such vast amounts of structured data allows for conceptualizing training ML models. PEACH [53], predicts channel quality based on images. If one has the CQI data for many UEs from TwinRAN and timestamped footage from a monitoring camera, one could train the PEACH model for their setup. Furthermore, one can adopt a continuous training approach by periodically training on a new dataset. Figure 6(a) shows the architecture in the cloud for deploying the use-case. The Azure DT instances produce events whenever a twin or a relationship is updated or created. These events are piped to an Azure Event Hub instance, a data streaming, processing, and ingestion platform. The processed events are streamed to the Azure Data Explorer instance, a data storage and analytics engine with an inbuilt persistent database. A Kusto Query Language (KQL) query from any platform can load data from the database to an ML model.
@@ VI-C Use Case III: What-if Scenarios
The DT in the cloud is the closest representation of the current digital state of the system. Network robustness and security can be greatly improved by conducting a what-if analysis by simulating different scenarios using a previous state of the network [39, 58, 59]. Figure 6(c) depicts the architecture for such a system using TwinRAN. A copy of the current TwinRAN state is made by simultaneously querying the entire instance and deploying it on many new instances. Azure Functions is a serverless computing service triggered by an event or an API call to make such one-to-many copes. The new instances of TwinRAN can then be connected to a simulation to conduct different what-if scenarios and compare the evolved twins. The vendor-agnostic capability of TwinRAN allows for such a seamless connection. We first populate a TwinRAN instance by the Flexric[60] with one real OpenAirInterface[61] gNB and 7 E2 agent emulators. The newly spawned copies of the TwinRAN instance were then connected to an NS-3 [47] simulation to test what-if scenarios. Some changes had to be made at the RIC-TwinRAN xApp connection depending on the three tested O-RAN RICs - Flexric, NS3-ORAN, and Near-RT RIC from the O-RAN software community.

### v2

@@ Abstract
The proliferation of 5G technology necessitates advanced network management strategies to ensure optimal performance and reliability. Digital Twin (DT)s have emerged as a promising paradigm for modeling and simulating complex systems like the 5G Radio Access Network (RAN). In this paper, we present TwinRAN, a DT of the 5G RAN built leveraging the Azure DT platform. TwinRAN is built on top of the Open RAN (O-RAN) architecture and is agnostic to the vendor of the underlying equipment. We demonstrate three applications using TwinRAN and evaluate the required resources and their performance for a network with  800 users and eight gNBs.
We first evaluate the performance and limitations of the Azure DT platform, measuring the latency under different conditions. The results from this evaluation allow us to optimize TwinRAN for the DT platform it uses. Then, we present the system’s architectural design, emphasizing its components and interactions. We propose that two types of twin graphs be simultaneously maintained on the cloud. The first is for intercell operations, keeping a broad overview of all the cells in the network. The second twin graph where each cell is spawned in a separate Azure DT instance for more granular operation and monitoring of intracell tasks. We evaluate the performance and operating costs of TwinRAN for each of the three applications. The TwinRAN DT in the cloud can keep track of its physical twin within a few hundred milliseconds, extending its utility to many 5G network management tasks - some of which are shown in this paper. The novel framework for building and maintaining a DT of the 5G RAN presented in this paper offers network operators enhanced capabilities, empowering efficient deployments and management.
@@ I Introduction
DTs, virtual replicas of physical systems, offer a promising approach to addressing these challenges [5]. Operators can simulate, monitor, and optimize network performance in real-time by creating a DT of the 5G RAN, enabling predictive management, efficient resource allocation, and rapid troubleshooting [6, 7]. DTs are being leveraged in many other industries, such as manufacturing, logistics, construction, and energy distribution [8]. At the forefront of the DT technology is the Microsoft Azure DT[9]. It is a Platform as a service (PaaS) to create, maintain, and interact with DTs on the cloud. The Azure cloud, where the DTs are hosted, is a production-grade cloud platform guaranteeing scalability, availability, and robustness. Moreover, other Azure cloud storage, analytics, and visualization services can be seamlessly connected to the DTs. Therefore, Azure DT presents a compelling opportunity for twinning the 5G RAN. Hence, we first measure the real-time performance of Azure DT under different conditions to make optimized design decisions for TwinRAN. The results presented from these measurements are a supplementary contribution to TwinRAN. To the best of our knowledge, this paper is the first to present a deep study of the behavior of Azure DT.
@@
The main contribution of this work is that it proposes a novel method and architecture for creating and managing a DT of the 5G RAN on the Azure cloud platform. Our approach leverages Azure DT’s standardized modeling, data integration, and ORAN’s non-invasive and vendor-agnostic properties. TwinRAN has three main features that underscore its practical utility: Non-Invasive: It does not require changes in the source code or hardware in the 5G RAN. By plugging the xApp on top of 5G RAN architectures, it builds seamlessly over the existing 5G components. Vendor Agnostic: It can work with all equipment and software vendors as long as they are O-RAN compliant. Multipurpose: The same DT can be simultaneously used for different applications as long as all the parameters required by the applications are modeled in the twins. Any application can connect to the DT to stream real-time updates or load historical data. Furthermore, this property extends TwinRANs’ utility for the future: any new application can be deployed without modifying the TwinRAN code. To present TwinRAN, it was first required to determine if the Azure DT platform suits this task. Secondly, for TwinRAN to be efficient, it was also necessary to characterize the behavior of Azure DT (especially in terms of latency) under different packet sizes and twin sizes. Thus, results from a thorough measurement study to ascertain the latency characteristics of Azure DT are a secondary contribution of our work. To the best of our knowledge, this paper is the first to present a deep study of the behavior of Azure DT.
The rest of the paper is organized as follows: the next section provides the background on Azure DTs and the O-RAN system. Following the background, we mention some related work and differentiate our contributions. In section IV, we evaluate the behavior of Azure DT. The learnings from this section are then used to present the architecture of TwinRAN in Section V. This section also elaborates on the Digital Twins Definition Language (DTDL) models and the xApp design. Section VI addresses the implementation of TwinRAN and introduces the three use cases, and Section VII evaluates the performance and required resources for TwinRAN, mainly focusing on pricing. Finally, in Section VIII, we conclude, presenting the limitations and point toward out future work.
@@ II-A4 Message Types
O-RAN defines several message types for communication between its components. Indication: Provides status updates and telemetry data from RAN elements to the Near-RT RIC. The two types of indication messages are report and insert. While the first gives periodic updates on subscription, the other is triggered asynchronously to notify the xApp of a specific event at the RAN. Policy: Convey policies from the Near-RT RIC to guide the behavior of RAN elements. The passing of such a policy function is reported to the TwinRAN instance by the TwinRAN xApp. Control: Used for configuring and managing RAN elements. While TwinRAN does not use control messages, we show an example in Section VI how they can be used in conjunction with TwinRAN to perform a handover.
@@ II-B1 Models
Models in Azure DT define the structure and properties of DTs. These models are created using the DTDL, specifying entities’ characteristics, behavior, and capabilities within the system. DTDL, is JSON-like and comprises a set of metamodel classes  [20]. Defining clear and consistent models ensures a standardized representation of the physical entities and their attributes [21, 19]. The main metamodel classes are: Interface: Defines the blueprint for a digital twin, outlining its structure, properties, and capabilities. Property: Describes attributes of a digital twin, holding specific static values. For example, the serial number of the thermometer. Relationship: Details the connections between digital twins, specifying how they interact and relate to each other—for example, the room and the thermometer and thermostat. Component: Represents a sub-entity of a digital twin by linking to an interface. It allows for modularity in defining models—for example, the window in a room. DTDL defines more classes, such as command and telemetry, which do not have a unique advantage in DTs [22] They are mainly useful for other Azure IoT services.
@@ III-B Digital Twin Platforms
Eclipse Ditto[28]An open-source framework part of the Eclipse IoT Working Group. It provides the tools to create, manage, and interact with digital twins. Similar to Azure DT, it has a JSON-based twin model defining language. However, twin instances are not maintained as knowledge graphs as in Azure DTs. It can be self-hosted and connected to various MQTT and HTTP protocol endpoints. MCX:  [44] An open-source model description and updating method via MQTT protocol. It connects to simulation frameworks such as MATLAB, Ansys, or custom Python classes. It facilitates connecting simulation frameworks with real-time data. AWS IoT TwinMaker[45] AWS IoT TwinMaker is a cloud-native platform rivaling the Azure DT platform in its capabilities and features. It connects to other AWS services for monitoring, analytics, and data ingestion. The models are defined in a JSON-like language, and the twin instances are maintained in a knowledge graph. Many web tools like Grafana dashboards are available for twin visualization and monitoring.
@@ IV Evaluating Azure DT
To ascertain whether any two distributions in our measurements are the same, we employ the Kolmogorov-Smirnov test with a p-value threshold of 0.05. The time of the day did not make any difference to any of the three values mentioned above. Similarly, updating or querying multiple twins on the same Azure DT instance did not increase either of the three values. When the model and update sizes were 100, we measured no difference between the first and last parameters’ lag. This indicates that the parameters on a twin are either not updated sequentially but rather parallelly, or the time to update all the parameters is very small compared to the overall lag and service time. The performance in terms of lag was highly scalable in terms of the number of twins. We did not notice any difference in the lag when whether there was one twin or 100 twins/instance. This is expected as the event route endpoint of each twin is different internally. Hence there is no pipeline blocking a twins update when parallelly updating a large number of twins at the same time. For $Y$ twins to be updated, it is like sending $Y$ different API calls.
@@ V-A Intercell and Intracell Instances
The limits placed by Azure DT provide us with considerable headroom on the message size. Let the maximum number of twins in an instance be $Y\in\mathbb{N}^{+}$. If every twin in the instance is updated once every $R\in\mathbb{R}^{+}$ seconds, then the Azure DT constraint is $R\geq 0.1$, and we get $Y=1000\cdot R$. 5G RAN systems are dynamic, and the maximum frame and slot length are 10 and 1 ms, respectively. Azure DT was initially built to represent the physical world using IoT, where the changes in the environment are slower than 5G. To maintain the fidelity of the DT in the cloud in TwinRAN, the update rate must be as high as possible within the limits placed by Azure DT. Updating all the twins at the maximum rate of $R=0.1$ seconds, we can have a maximum of 100 twins per instance. Any relationship can be defined between two twins in the same instance. Thus, a problem arises in organizing the twins to achieve scalability while keeping to 100 twins per instance limitation and keeping all the relationships between entities. At the same time, the models’ size must be kept small to decrease the lag of the twins. The models must also be organized to fit the entities traditionally used in cellular networks. Hence, we propose to split TwinRAN into two instance types.
From the perspective of a 5G RAN operator, the tasks can be divided into two main categories: intracell and intercell. Intracell operation tasks include efficient scheduling of users based on their required Quality of Service (QoS), minimizing operating costs, network security, etc. Intercell operations are handover management, interference mitigation, etc. Therefore, we propose two twin-graphs simultaneously maintained in the cloud: TwinRAN-cell and TwinRAN-multi. TwinRAN-multi keeps a broad overview of the entire network. Each cell is defined in one DTDL model and neighboring cells are connected with relationships. Each UE is one property of the gNB, with SINR as its value. TwinRAN-cell spawns a new instance for every gNB. Here, the gNB and UE are two different DTDL models with a relationship between them if the UE is connected. We present the DTDL models in detail in the following subsection. A large model of a cell with detailed models of the gNB and UE as components would allow us to simultaneously manage 100 cells, each updating 10 times a second. However, a cell containing 100s of UEs would make the update sizes large and increase the twins’ lag, service time, and query time. Moreover, this would impact the system’s scalability as an additional parameter in the UE model and severely affect the overall lag. We can keep these values low by employing a disaggregated approach that simultaneously maintains TwinRAN-cell and TwinRAN-multi instances.
@@ V-B DTDL Models
As mentioned in Section II-B, models defined by DTDL provide a structure to each twin. TwinRAN defines three models: the cell, User Equipment (UE), and the gNB. The cell is used in TwinRAN-multi. The input to our modeling step is the information we can obtain via different O-RAN messages mentioned in Section II-A. The information is divided into Information Elements (IE)s, which are structured data units used within the O-RAN protocol messages to convey specific parameters. While the amount of information is substantial, and thus, one can make very detailed models, we make the smallest possible model useful for all the three use cases mentioned in Section VI.
Table I shows the two DTDL interfaces for the TwinRAN-multi instance. The gNB is already initialized with the maximum number of UEs it can accommodate (99), regardless of their connection. This is because changing a model in the live execution environment requires more than 1 second, and hence, we do not need to change the model for every handover. Neighboring gNBs are connected with a relationship whose value is the received power from that gNB. The properties of the UE that are important to demonstrate our use cases are its location and received power.
@@
Tx Power float 20-50 DL CQI int 0-15
Relationship int UE UL MCS int 0-28
Connected UEs int 0-99 UL MCS int 0-28
@@ V-C TwinRAN xApp
The TwinRAN xApp updates the DT on the cloud and manages the instances, creating twins and maintaining relationships. The information obtained from the E2 Node is provided by different RAN functions [18]. TwinRAN subscribes to the required RAN functions via the Near-RT RIC as shown in Figure 1. For periodic indication messages(called as report, see Sec.  II-A), the subscription request also contains a parameter called granularity. This parameter represents the frequency at which the report messages should be sent by the E2 Node. The gNB in the E2 setup request provides a cell ID and PLMN information used to spawn a new cell in the TwinRAN-muli instance. Meanwhile, another instance is created in the Azure DT cloud, where the gNB twin is created. One expects this to happen rarely, as cells do not frequently appear and disappear from the operator’s network.
@@ VI-A Use Case I: Data Warehouse for Machine Learning
A data warehouse is a centralized repository that stores large volumes of structured data from multiple sources. Access to such vast amounts of structured data allows for conceptualizing training ML models. PEACH [53] predicts channel quality based on images. If one has the CQI data for many UEs from TwinRAN and timestamped footage from a monitoring camera, one could train the PEACH model for their setup. Furthermore, one can adopt a continuous training approach by periodically training on a new dataset. Figure 6(a) shows the architecture in the cloud for deploying the use-case. The Azure DT instances produce events whenever a twin or a relationship is updated or created. These events are piped to an Azure Event Hub instance, a data streaming, processing, and ingestion platform. The processed events are streamed to the Azure Data Explorer instance, a data storage and analytics engine with an inbuilt persistent database. A Kusto Query Language (KQL) query from any platform can load data from the database to an ML model.
@@ VI-C Use Case III: What-if Scenarios
The DT in the cloud is the closest representation of the current digital state of the system. Network robustness and security can be greatly improved by conducting a what-if analysis by simulating different scenarios using a previous state of the network [39, 58, 59]. Figure 6(c) depicts the architecture for such a system using TwinRAN. A copy of the current TwinRAN state is made by simultaneously querying the entire instance and deploying it on many new instances. Azure Functions is a serverless computing service triggered by an event or an API call to make such one-to-many copies. The new instances of TwinRAN can then be connected to a simulation to conduct different what-if scenarios and compare the evolved twins. The vendor-agnostic capability of TwinRAN allows for such a seamless connection. We first populate a TwinRAN instance by the Flexric[60] with one real OpenAirInterface[61] gNB and 7 E2 agent emulators. The newly spawned copies of the TwinRAN instance were then connected to an NS-3 [47] simulation to test what-if scenarios  [^1] Some changes had to be made at the RIC-TwinRAN xApp interface depending on the three tested O-RAN RICs - Flexric, NS3-ORAN, and Near-RT RIC from the O-RAN software community. This is a limitation because true interoperability is still not achieved by any ORAN implementation or vendor and until this original goal of ORAN is achieved, small vendor-specific changes will need to be made at all points in the data chain. .